OpenStudy (anonymous):

Help please!! find dy/dx by implicit differentiation. cot(y)=x-6y i got x/(6-(csc(y)^2) and is wrong

OpenStudy (anonymous):

i got your back

OpenStudy (anonymous):

so clam down ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so can you give me more details

OpenStudy (anonymous):

i need more details ok

OpenStudy (dumbcow):

looks pretty good actually, except you forgot to take derivative of the x \[\frac{dy}{dx} = \frac{1}{6-\csc^2 y}\]

OpenStudy (anonymous):

that's the question... the answer i got by doing, i just followed an example that's in my book, im not sure if its right: -(csc y)^2 dy=x-6y and then i think i solved for dy and it got dy/dx=x/(6-(csc y)^2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@dumbcow thanks, i see my mistake now, thanks